What is the pH of a 0.485 M ammonia solution?

Ammonia, rm NH_3, is a weak base with a K_{rm b} value of 1.8times 10^{-5}.

a)What is the pH of a 0.485 it Mammonia solution?

Express your answer numerically to two decimal places.

b)What is the percent ionization of ammonia at this concentration?

Express the percentage numerically to three significant figures.Assignment 29 Ch.16
Answer #1

Given Kb = 1.8x10-5

1.8x10-5 = x2 / 0.485-x
x = [OH-] = 0.0029 M
pOH = -log (OH-)= 2.53
pH = 14 - 2.53 = 11.47

answered by: Special